What will be the temperature reading in fahrenheit scale when the reading is 310k? 

Grade:7

1 Answers

Arun
3 years ago
Dear student
 
Kelvin = 310
 
Hence on centigrade scale = 310 -273 = 37
 
Now
Use formula
 
C/5 = F -32/9
 
37/5 = F-32/9
 
66.6 = F -32
 
F = 32+66.6 = 98.6
 
Hope you understand